### v1

Abstract
Spatial heterogeneity widely exists in many applications, such as in ocean science, where the temperature-salinity (T-S) relationship in thermohaline circulation varies across different geographical locations and depths. While spatial regression models are powerful tools for this purpose, they often face challenges in simultaneously estimating spatial parameters, detecting heterogeneity boundaries, and performing adaptive modeling, especially in complex systems. This paper proposes a Fused Lasso regression model with an Adaptive minimum spanning Tree (FLAT) to address these challenges in a unified framework. Specifically, FLAT constructs an adaptive minimum spanning tree guided by both spatial proximity and coefficient dissimilarity, and incorporates a spatial heterogeneity penalty to capture the underlying structure. A subsequent spatial clustering algorithm then identifies discrete heterogeneity boundaries, such as oceanic thermohaline fronts. Numerical simulations confirm that FLAT significantly outperforms classic spatial regression models in both coefficient estimation and heterogeneity detection. An empirical analysis with Atlantic Ocean data further demonstrates FLAT’s capability to elucidate region-specific thermohaline compensation mechanisms and to detect surfaces with inverse T-S relationships. These findings advance the mechanistic understanding of T-S compensation dynamics in the Antarctic Intermediate Water region.
keywords:
Spatial heterogeneity , Fused regularization , Minimum spanning tree , Thermohaline circulation , Temperature-salinity relationship
11footnotetext:   All authors equally contribute.22footnotetext: * Corresponding author. Email: zxing@nd.edu
1 Introduction
Spatial heterogeneity is a fundamental feature of geophysical processes in the modern earth system, profoundly affecting phenomena from urban heat islands to the oceanic thermohaline circulation (THC) (Anselin, 2013; Talley, 2011). The THC, a global conveyor belt driven by density gradients from temperature and salinity variations, plays a pivotal role in regulating Earth’s climate (Clark et al., 2002). Its meridional overturning component transports vast amounts of heat poleward and is intrinsically linked to critical diapycnal mixing processes (Talley, 2013). A key challenge in oceanography is that this system exhibits a dual nature of spatial non-stationarity: it features both smooth, continuous variations within water masses and sharp, discrete transitions at their boundaries (Fotheringham et al., 2017; Ferrari et al., 2001). Accurately capturing this duality is crucial for understanding oceanic dynamics, yet it poses a significant challenge for existing analytical methods.
On one hand, spatial systems exhibit continuous heterogeneity, where spatial variables and their relationships vary smoothly across geographic space (Bailey et al., 1995). This phenomenon, also known as spatial continuous non-stationarity (Fotheringham et al., 2017), is evident in numerous environmental contexts. For instance, in climatology, temperature tends to decline and precipitation increases with elevation due to reduced atmospheric pressure and diminished heat absorption capacity. Similarly, in terrestrial hydrology, soil moisture-temperature relationships often vary continuously across space (Zhang et al., 2024). However, focusing solely on this smooth aspect of reality provides an incomplete picture, as many systems are also structured by abrupt, structural divisions. On the other hand, these abrupt divisions manifest as spatially discrete heterogeneity, referring to sharp transitions or categorical differences in variable relationships across space (Ferrari et al., 2001; Anselin, 2010). Such discontinuities are commonly observed in dynamic systems, from sudden shifts in rainfall regimes along mountain ridges (Luo et al., 2021) to sharp thermohaline interfaces marked by steep temperature gradients in the ocean (Li and Yang, 2021). Therefore, The fundamental scientific challenge lies in modeling both continuous and discrete patterns that coexist within the same geographic space.
To model such heterogeneity, several spatial regression models and their extensions have been developed (LeSage and Pace, 2009; Ward and Gleditsch, 2018). Classic approaches include the spatial lag model (Lam and Souza, 2020), the spatial Durbin model (Anselin, 2013), and geographically weighted regression (GWR), which captures local spatial dependencies (Brunsdon et al., 1996). However, these models are primarily designed to estimate spatially varying coefficients, thereby capturing only the continuous aspect of heterogeneity. They generally do not aim to detect abrupt changes or spatial boundaries in the coefficient surfaces that are of central importance in fields like oceanography, such as the inverse temperature-salinity surface in thermohaline circulation. To address such cases, some empirical studies employ two-step procedures that involve space partitioning (Chen et al., 2024; Duque et al., 2011) and clustering-based segmentation techniques (Raudenbush, 2002; Van Dusen and Nissen, 2019). However, classic spatial models (Ward and Gleditsch, 2018) often fail to capture sharp transitions stably at regional boundaries. Since real-world geographic spaces often exhibit both continuous and discrete forms of spatial heterogeneity, it is crucial to develop adaptive modeling methods with capability to capture local smoothness while also identifying abrupt spatial boundaries of coefficients.
We propose a novel spatial regression framework, the Fused Lasso with an Adaptive minimum spanning Tree (FLAT), which addresses the challenges by introducing a structured adaptive regularization that captures both spatial heterogeneity in coefficients and sparsity in high-dimensional features. Specifically, we construct fused penalization terms guided by minimum spanning tree structures that incorporate both spatial proximity and coefficient similarity, and we incorporate adaptive weights to enhance estimation accuracy and computational efficiency. In addition, we include extra penalization terms to achieve simultaneous variable selection. A density-based spatial clustering method can be conducted to further detect the non-convex spatial regions with heterogeneity in applications, such as the detection of inverse temperature-salinity surface in the ocean. Comprehensive simulations and a real-world application to thermohaline circulation data demonstrate that FLAT yields accurate estimation, robust feature selection, and can provide additional unique insights in empirical analysis with interpretable results, which is applicable for data-driven discovery in geophysical sciences and other interdisciplinary applications.
The rest of the article is organized as follows. Section 2 introduces the methodology of the fused Lasso regression with adaptive minimum spanning tree. Simulation studies are provided in Section 3, and a real application of our method on thermohaline circulation is shown in Section 4. Section 5 concludes this article with a discussion. Technical details and additional details are included in the Appendix.
2 Fused Lasso Regression With Adaptive Minimum Spanning Tree
To characterize both the complex relationships and heterogeneity in spatial data, we denote the spatial data at the $i$-th spatial location $s_{i}$ as $\{\bm{x}(s_{i}),y(s_{i})\}$ for $i\in[n]$, where $\bm{x}(s_{i})=\left[x_{1}(s_{i}),\cdots,x_{p}(s_{i})\right]^{\top}$ is the $p$-dimensional vector, $y(s_{i})$ is the corresponding response and $[n]=\{1,\cdots,n\}$. Then we consider a spatially varying coefficient regression model as below:
$y(s_{i})=\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})+\epsilon(s_{i}),$ (1)
[…]
We firstly generate the minimum spanning tree $\mathbb{E}_{0}$ based on spatial distances and initialize the parameter estimators $\hat{\bm{\beta}}^{0}(s_{i}),i\in[n]$ by optimizing the following target function
$\mathop{\arg\min}\limits_{\bm{\beta}}\frac{1}{n}\sum\limits_{i=1}^{n}\left(y(s_{i})-\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})\right)^{2}+\lambda\sum\limits_{k=1}^{p}\sum\limits_{(s_{i},s_{j})\in\mathbb{E}_{0}}|\beta_{k}(s_{i})-\beta_{k}(s_{j})|.$ (2)
Then we generate the full connection graph among spatial locations based on the $L_{2}$ distance between coefficients as
$\tilde{d}(s_{i},s_{j})=\|\hat{\bm{\beta}}^{0}(s_{i})-\hat{\bm{\beta}}^{0}(s_{j})\|_{2},$ (3)
and obtain the spatial linkage structure $\mathbb{E}$ via Prim’s algorithm based on the estimated distance matrix.
To introduce discrete spatial heterogeneity in the spatial regression model for continuous spatial heterogeneity, we construct the connectivity matrix $\mathbf{H}\in\mathbb{R}^{(n-1)\times n}$ from $\mathbb{E}$, which represents the minimum spanning tree formed by $(n-1)$ edges and $n$ vertices. The elements $\mathbf{H}_{l,i}$ of $\mathbf{H}$ are defined based on the connection status of vertices $i(1,...,n)$ along the edges $l(1,...,n-1)$. If the $l$-th edge $(s_{i},s_{j})\in\mathbb{E}$, then $\mathbf{H}_{l,i}=1$, $\mathbf{H}_{l,j}=-1$, and $\mathbf{H}_{l,v}=0$ for $v\neq i,j$. Thus, we have
$\|\mathbf{H}\bm{\beta}_{k}\|_{1}=\sum_{(s_{i},s_{j})\in\mathbb{E}}|\bm{\beta}_{k}(s_{i})-\bm{\beta}_{k}(s_{j})|,$ (4)
where $\bm{\beta}_{k}=(\beta_{k}(s_{1}),...,\beta_{k}(s_{n}))^{\top},\quad k=1,...,p$ represent the coefficient values of the $k$-th explanatory variable at each observation point. Based on equation (4), $\|\mathbf{H}\bm{\beta}_{k}\|_{1}$ represents the heterogeneity penalty based on the adaptive minimum spanning tree for the $k$-th explanatory variable.
To further improve the finite sample performance of both estimations and detection of discrete spatial heterogeneity, we choose the adaptive weight between any two spatial $(s_{i},s_{j})$ as $\pi_{(s_{i},s_{j})}=1/\tilde{d}(s_{i},s_{j})^{\gamma}$ for all $(s_{i},s_{j})\in\mathbb{E}$ and 00 otherwise. We follow the recommendation of Zou (2006) and set $\gamma=1$ , where $\gamma$ is a constant adjusting the adaptiveness of the penalty. Intuitively, spatial locations with small distances in the initial coefficients, denoted by $\tilde{d}(s_{i},s_{j})$, are more likely to belong to the same region or share similar underlying processes. In this context, larger adaptive weights impose stronger penalization on discrepancies between nearby coefficients, thereby encouraging local homogeneity and smoothing. This leads to estimated coefficients that are more likely to align in regions of weak heterogeneity, effectively capturing both smooth transitions and abrupt changes. Based on this intuition, the adaptive penalization on spatial heterogeneity is further formulated as:
$P_{\lambda_{1}}(\bm{\beta})$ $=\sum_{(s_{i},s_{j})\in\mathbb{E}}P_{\lambda_{1}}\left(\bm{\beta}(s_{i})-\bm{\beta}(s_{j})\right)$
$=\lambda_{1}\sum_{k=1}^{p}\sum_{i=1}^{n}\sum_{j\in\mathcal{N}_{i}}\pi_{(s_{i},s_{j})}\left|\beta_{k}(s_{i})-\beta_{k}(s_{j})\right|,$ (5)
where $\mathbb{E}$ represents the edge set derived from the minimum spanning tree based on the correlation similarity matrix, and $\mathcal{N}_{i}$ denotes the neighboring nodes of node $i$ in the minimum spanning tree. Ultimately, the penalty for spatial heterogeneity can be rewritten as:
$P_{\lambda_{1}}(\bm{\beta})=\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1},$ (6)
where the weight matrix $\bm{\pi}\in\mathbb{R}^{(n-1)\times(n-1)}$ represents the weight of the edge $(s_{i},s_{j})$.
Regarding the sparsity of high-dimensional spatial features, we conduct variable selection via additional Lasso penalty to shrink coefficients to zero and to decrease model complexity. The objective function of FLAT is
$L_{\text{FLAT}}(\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\left(y(s_{i})-\bm{\beta}^{\top}(s_{i})\mathbf{x}(s_{i})\right)^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are tuning parameters controlling the strength of penalization to capture the discrete spatial heterogeneity and the sparsity in spatial features. The adaptive weights $\bm{\pi}$ further enhance the detection of spatial heterogeneity while reducing the estimation error for the spatial locations with homogeneity.
For simplicity, we use matrix notation and rewrite the objective function of FLAT (7) as
$L_{\text{FLAT}}(\bm{\beta})=\frac{1}{2}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\pi}\mathbf{H}\bm{\beta}_{k}\|_{1}+\lambda_{2}\sum_{k=1}^{p}\|\bm{\beta}_{k}\|_{1},$ (8)
where $\mathbf{y}=[y(s_{1}),...,y(s_{n})]^{\top}\in\mathbb{R}^{n}$ represents the vector of response values for $n$ locations. $\mathbf{X}\triangleq[\text{diag}(\mathbf{x}_{1}),...,\text{diag}(\mathbf{x}_{p})]\in\mathbb{R}^{n\times np}$, where $\mathbf{x}_{k}$ represents the $k$-th covariate vector of length $n$, and the observation vector is $\mathbf{x}_{k}\triangleq(x_{k}(s_{1}),...,x_{k}(s_{n}))^{\top}\in\mathbb{R}^{n}$ is the corresponding design matrix with spatial feature. We denote $\bm{\beta}=[\bm{\beta}^{\top}(s_{1}),...,\bm{\beta}^{\top}(s_{n})]^{\top}\in\mathbb{R}^{np}$, where $\bm{\beta}(s_{i})\triangleq[\beta_{1}(s_{i}),...,\beta_{p}(s_{i})]^{\top}\in\mathbb{R}^{p}$.
To further simplify the notation, we define the matrix $\mathbf{\tilde{H}}$ and reparameterize coefficient matrix accordingly
$\mathbf{\tilde{H}}\triangleq\begin{bmatrix}\pi\mathbf{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\mathbf{1}^{\top}\end{bmatrix}\in\mathbb{R}^{n\times n},\leavevmode\nobreak\ \leavevmode\nobreak\ \bm{\theta}_{k}=\begin{bmatrix}\pi\mathbf{H}\\ \frac{\lambda_{2}}{\lambda_{1}}\mathbf{1}^{\top}\end{bmatrix}\bm{\beta}_{k}=\mathbf{\tilde{H}}\bm{\beta}_{k},$ (9)
where $\bm{\theta}_{k}\triangleq[\theta_{k}(s_{1}),...,\theta_{k}(s_{n})]^{\top}\in\mathbb{R}^{n}$. Since $\mathbf{\tilde{H}}$ is invertible, we reparameterize $\bm{\beta}$ as equation (9) and express $\bm{\beta}_{k}$ as $\bm{\beta}_{k}=\mathbf{\tilde{H}}^{-1}\bm{\theta}_{k}$. Based on this transformation, we reparameterize the design matrix as
$\mathbf{\tilde{X}}\triangleq\left[\text{diag}(\mathbf{x}_{1})\mathbf{\tilde{H}}^{-1},...,\text{diag}(\mathbf{x}_{p})\mathbf{\tilde{H}}^{-1}\right].$ (10)
With the reparameterization defined above, we rewrite the object function of proposed FLAT approach in (7) as
$L_{\text{FLAT}}(\bm{\theta})=\frac{1}{2}\|\mathbf{y}-\mathbf{\tilde{X}}\bm{\theta}\|_{2}^{2}+\lambda_{1}\sum_{k=1}^{p}\|\bm{\theta}_{k}\|_{1},$ (11)
where $\bm{\theta}\triangleq[\bm{\theta}_{1}^{\top},...,\bm{\theta}_{p}^{\top}]^{\top}\in\mathbb{R}^{np\times 1}$. We define the FLAT estimator of the coefficients as $\hat{\bm{\theta}}_{\mathrm{FLAT}}=\arg\min_{\bm{\theta}\in\Theta}L_{\text{FLAT}}(\bm{\theta})$. We can obtain the FLAT estimator of original regression coefficients using $\hat{\bm{\beta}}_{\mathrm{FLAT}}=\mathbf{\tilde{H}}^{-1}\hat{\bm{\theta}}_{\mathrm{FLAT}}$.
By minimizing (11), $\hat{\bm{\theta}}_{\mathrm{FLAT}}$ can detect the discrete spatial heterogeneity in effects while estimating the continuous spatial heterogeneity, and also select high-dimensional spatial features. To indicate non-convex spatial clusters in terms of feature effects, some classic density-based clustering algorithm (Ester et al., 1996; Miller and Han, 2009) can be used to further obtain the spatial surface in specific applications, such as the inverse temperature-salinity surface in thermohaline circulation. The details of the FLAT algorithms are summarized in Algorithm 1.
1:  Input: Spatial data $\{(\bm{x}(s_{i}),y(s_{i}))\}_{i=1}^{n}$, parameters $\lambda_{1},\lambda_{2},\gamma$
2:  Initialization of pairwise distances via (3) and obtain the adaptive weights $\pi_{(s_{i},s_{j})}$ in (2)
3:  Obtain the fused structure $\bm{H}$ in (4) based on edge set $\mathbb{E}$, which is built by minimum spanning tree from Prim’s algorithm
4:  Reparameterization defined in (9) Compute adaptive weights $\bm{\pi}$ in (8)
5:  Formulate objective function and conduct reparameterization via (11)
6:  repeat
7:     for all  $j=1$ to $np$ do
8:        Update via soft-thresholding operator $\mathcal{S}_{\lambda}(x)=\text{sign}(x)(|x|-\lambda)_{+}$ and
$\theta_{j}^{(t+1)}=\mathcal{S}_{\lambda_{1}}\left(\theta_{j}^{(t)}-\eta\sum_{i=1}^{n}\tilde{X}_{ij}(y_{i}-\hat{y}_{i}^{(t)}+\tilde{X}_{ij}\theta_{j}^{(t)})\right)$
9:     end for
10:  until $\|\bm{\theta}^{(t+1)}-\bm{\theta}^{(t)}\|_{2}<10^{-6}$
11:  Obtain $\hat{\bm{\theta}}_{\mathrm{FLAT}}$ and the estimated coefficients $\hat{\bm{\beta}}_{\mathrm{FLAT}}$
12:  Initialization of the hyper-parameter sequences $\epsilon$ and $minPts$, set $cluster.id=0$
13:  for all $(\epsilon,minPts)$ do
14:     Initialize all locations as unvisited
15:     for all unvisited locations $s_{i}$ do
16:        Obtain its $\epsilon$-neighborhood $N_{\epsilon}(s_{i})=\{s_{j}:\|\hat{\bm{\beta}}_{\mathrm{FLAT}}(s_{j})-\hat{\bm{\beta}}_{\mathrm{FLAT}}(s_{i})\|_{2}\leq\epsilon\}$
17:        if $|N_{\epsilon}(s_{i})|\geq minPts$ then
18:           Create new cluster: $cluster.id\leftarrow cluster.id+1$
19:           for all $s_{j}\in N_{\epsilon}(s_{i})$ do
20:              Compute $N_{\epsilon}(s_{j})$
21:              if $|N_{\epsilon}(s_{j})|\geq minPts$ then
22:                 Merge $N_{\epsilon}(s_{j})$ into $N_{\epsilon}(s_{i})$
23:              else
24:                 Mark $s_{j}$ as noise
25:              end if
26:           end for
27:           Expand cluster: Add $N_{\epsilon}(s_{i})$ to cluster $C_{cluster.id}$
28:           Mark all locations in $C_{cluster.id}$ as visited
29:        end if
30:     end for
31:     Evaluate and select the clustering labels
32:  end for
33:  Output: Estimated coefficients $\hat{\bm{\beta}}_{\mathrm{FLAT}}$, detected spatial clusters
Algorithm 1 FLAT Algorithm with Adaptive MST
3 Simulation
3.1 Simulation settings
To fully characterize marine conditions where the response variables of regression models often exhibit distinct spatial structures in geoscientific research, we consider two-dimensional spatial data in the simulations, which is commonly considered in marine across longitude and latitude coordinates or longitude and depth. Without losing generality, we assume a $[0,1]\times[0,1]$ spatial domain with $n=1,000$ randomly distributed observation points $s_{1},\dots,s_{n}\in\mathbb{R}^{2}$.
In all settings of our simulations, the responses at spatial locations $\{s_{1},\dots,s_{n}\}$ are generated based on spatial regression model
$y(s_{i})=\beta_{1}(s_{i})x_{1}(s_{i})+\beta_{2}(s_{i})x_{2}(s_{i})+\beta_{3}(s_{i})+\epsilon(s_{i}),\quad i=1,\dots,n,$ (12)
where $\epsilon(s_{i})$ represent the independent normal error terms with zero mean and standard deviation $\sigma=0.1$. We use different approaches to estimate the parameters $\{\beta_{1}(s_{i}),\beta_{2}(s_{i}),\beta_{3}(s_{i})\}_{i\in[n]}$ based on the generated spatial data $\{\bm{x}(s_{i}),y(s_{i})\}_{i\in[n]}$ and compare their performance in capturing both continuous spatial heterogeneity and discrete spatial heterogeneity.
The spatial dependence among spatial covariates is specially designed with different covariance structure. In all simulations, two independent Gaussian stochastic processes $\{z_{1}(s_{i})\}_{i=1}^{n}$ and $\{z_{2}(s_{i})\}_{i=1}^{n}$ are generated with zero mean and designed covariances $\text{cov}(z_{k}(s_{i}),z_{k}(s_{j}))=\exp(-\|s_{i}-s_{j}\|_{2}/\phi),\ k=1,2.$
In the generating process, parameter $\phi$ controls the spatial heterogeneity. Then the spatial features $x_{1}(s_{i})$ and $x_{2}(s_{i})$ are generated from $z_{1}(s_{i})$ and $z_{2}(s_{i})$ by linear transformation. Specifically, we set $x_{1}(s_{i})=z_{1}(s_{i})$ and $x_{2}(s_{i})=rz_{1}(s_{i})+\sqrt{1-r^{2}}z_{2}(s_{i})$, where the parameter $r$ controls the level of collinearity. We set $r=0.75$ for a strong correlation design and generate five datasets with $\phi=0.2,0.4,0.6,0.8,1$ for different levels of spatial correlation, which are labeled as “DGP1”, $\cdots$, “DGP5”, respectively. The true values of the parameters $\beta_{1},\beta_{2},\beta_{3}$ are set as shown in Figure 1(a)-(c) respectively, where different clustering patterns are adopted for three different parameters.
To quantify the performance of estimations for dimension $k\in[p]$ and location $i\in[n]$, we focus on the three measures with $T=100$ independent replications: (1) Root Mean Squared Error (RMSE), which is defined as $\text{RMSE}_{k}^{(i)}=\left[\sum\limits_{t=1}^{T}\left(\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})\right)^{2}/T\right]^{1/2}$ (2) Mean Absolute Error (MAE), which is defined as $\text{MAE}_{k}^{(i)}=\sum\limits_{t=1}^{T}|\hat{\beta}_{k}^{t}(s_{i})-\beta_{k}(s_{i})|/T$; (3) Standard Deviation (SD), which is defined as $\text{SD}_{k}^{(i)}=\left[\sum\limits_{t=1}^{T}\left(\hat{\beta}_{k}^{t}(s_{i})-\bar{\hat{\beta}}_{k}(s_{i})\right)^{2}/T\right]^{1/2}$, where $\bar{\hat{\beta}}_{k}(s_{i})=\sum\limits_{t=1}^{T}\hat{\beta}_{k}^{t}(s_{i})/T$. Smaller values in RMSE, MAE and SD indicate more accurate and stable estimations for parameters in model (12). As we have in total $np$ parameters to estimate, the mean values of the aforementioned measures across all spatial points $i\in[n]$ for each dimension $k\in[p]$ are computed and reported in Table 1.
To assess the performance in detecting discrete spatial heterogeneity in each dimension $k\in[p]$, we consider four widely used measures for clustering: Rand Index (RI), Adjusted Rand Index (ARI), Silhouette Coefficient (SC) and Calinski-Harabasz Index (CHI). Higher values of RI, ARI, SC and CHI reflect a greater consistency between the detected clusters and the true group labels. Detailed definitions and further discussion of these metrics are included in A.
We compare our FLAT method with two classic works: Geographically Weighted Regression (GWR) model (Brunsdon et al., 1996) and Spatially Clustered Coefficient (SCC) model (Li and Sang, 2019). To mitigate the impact of errors due to randomness, we conduct $T=100$ repeats for each setting and evaluate the estimating and clustering performance of different methods.
3.2 Simulation results
We firstly illustrate the finite sample performance of our method FLAT through scatter plots in Figure 1 and provide a direct visual comparison of estimated coefficients by different methods. The first row with three plots (a)-(c) shows the true parameters in data-generating process, and other rows show the estimated parameters by GWR, SCC and proposed FLAT, respectively. The three columns represent the spatial coefficients $\beta_{1}(s_{i})$, $\beta_{2}(s_{i})$ and $\beta_{3}(s_{i})$, and the color represents the quantities of the parameters. From the sub-figure (a)-(c), there’s different layers in each scatter plot, which indicates the cluster patters in different dimension. Sub-figure (d)-(f) shows that GWR can hardly detect the heterogeneity in spatial parameters, and FLAT provides more accurate estimations in sub-figure (j)-(l) compared to GWR. While the SCC can detect part of the spatial heterogeneity, FLAT excels in boundary identifications as marked in the boxed areas of sub-figure (g) and (h).
Table 1: The comparison of the performance in estimation with 5 data-generating processes
Evaluation Parameter Method DGP1 DGP2 DGP3 DGP4 DGP5
RMSE $\beta_{1}$ GWR 3.0521 4.2638 5.1742 5.9329 6.7977
SCC 0.8612 1.0648 1.2451 1.3882 1.5690
FLAT 0.8090 0.9800 1.1503 1.2817 1.4446
$\beta_{2}$ GWR 3.0344 4.2502 5.1628 5.9215 6.8203
SCC 0.9390 1.1372 1.3132 1.4588 1.6872
FLAT 0.8894 1.0454 1.2160 1.3582 1.5665
$\beta_{3}$ GWR 3.0311 4.5394 5.6082 6.4590 7.5956
SCC 0.9015 1.1835 1.4219 1.5988 1.8425
FLAT 0.8502 1.0845 1.3177 1.4848 1.7280
MAE $\beta_{1}$ GWR 2.3340 3.2390 3.9096 4.4751 5.0945
SCC 0.5856 0.7330 0.8726 0.9953 1.1371
FLAT 0.5370 0.6645 0.7906 0.8984 1.0293
$\beta_{2}$ GWR 2.3247 3.2121 3.8829 4.4486 5.0979
SCC 0.6520 0.8019 0.9443 1.0641 1.2616
FLAT 0.5940 0.7191 0.8520 0.9684 1.1459
$\beta_{3}$ GWR 2.0214 2.8968 3.5364 4.0649 4.7283
SCC 0.5698 0.7448 0.8997 1.0234 1.2073
FLAT 0.5395 0.6884 0.8280 0.9445 1.1196
SD $\beta_{1}$ GWR 2.9521 4.1770 5.0925 5.8530 6.7234
SCC 0.7213 1.4430 1.0591 1.1813 1.3341
FLAT 0.6888 0.8401 0.9865 1.0991 1.2338
$\beta_{2}$ GWR 2.9500 4.1666 5.0833 5.8433 6.7586
SCC 0.7721 1.7088 1.0781 1.1950 1.3522
FLAT 0.7501 0.8764 1.0102 1.1270 1.2729
$\beta_{3}$ GWR 2.7790 4.3506 5.4442 6.3084 7.4337
SCC 0.7644 1.8384 1.2170 1.3703 1.5622
FLAT 0.7151 0.9165 1.1237 1.2681 1.4605
Figure 1: True spatial parameters $\beta_{k}$ and estimated parameters $\hat{\beta}_{k}$ for $k=1,2,3$ by GWR, SCC and FLAT: In each subgraph, the coordinate axes correspond to two spatial dimensions and the points represent the spatial locations $s_{i}$ accordingly. The dataset is generated from DGP5 and the values of parameters and estimations are indicated by color.
The comparison based on the measures of estimations shows the similar conclusion in Table 1, where the mean values of the different measures across all spatial points $i\in[n]$ for each dimension $k\in[p]$ are computed and reported for each individual parameter $\beta_{1}$, $\beta_{2}$ and $\beta_{3}$ under the different data-generating settings. For each dimension of spatial parameter $\bm{\beta}$, FLAT achieves the dominated performance in estimation in terms of RMSE, MAE and SD compared to the classic methods GWR and SCC. Besides, FLAT shows more stable estimations with increasing local dependency from DGP1 to DGP5, while GWR can hardly handle the strong spatial correlation.
The comparison based on the measures of clustering is shown in Table 2. As shown in the first row of Figure 1, true spatial parameters have spatial clustering patterns. We conduct the identical standard DBSCAN clustering algorithm with the estimations by GWR, SCC and FLAT and calculate the RI, ARI, SC, and CHI, which measure the similarity between the discrete spatial heterogeneity and the detected clusters of coefficients in different dimensions. Thus, more accurate clustering results will indicate better capture of discrete spatial heterogeneity. As the Table 2, FLAT achieves better performance with higher RI, ARI, SC and CHI compared to the classic methods GWR and SCC, which indicates the effectiveness of FLAT in detecting the discrete spatial heterogeneity in practice. Similar conclusions are found with other data-generating processes, and we include the additional numerical results in B.
Table 2: Evaluations based on clustering results of estimated parameters $\hat{\bm{\beta}}$ by different methods
Data-generating process Parameter Method RI ARI SC CHI(×$\times 10^{4}$)
DGP1 $\beta_{1}$ GWR 0.6411 0.2453 0.4537 0.2041
SCC 0.9117 0.7898 0.8685 2.3816
FLAT 0.9180 0.8051 0.8791 3.3122
$\beta_{2}$ GWR 0.7446 0.2769 0.4163 0.5611
SCC 0.9185 0.7490 0.8378 3.1291
FLAT 0.9289 0.7804 0.8577 4.0711
$\beta_{3}$ GWR 0.8084 0.3576 0.5163 0.9952
SCC 0.9346 0.7709 0.8713 4.3364
FLAT 0.9370 0.7818 0.8835 4.7318
DGP5 $\beta_{1}$ GWR 0.6101 0.1260 0.1879 0.3858
SCC 0.8056 0.5650 0.7494 0.7449
FLAT 0.8227 0.5951 0.7680 1.0421
$\beta_{2}$ GWR 0.6760 0.1487 0.1363 0.7457
SCC 0.8307 0.5023 0.7572 1.6862
FLAT 0.8471 0.5386 0.7754 2.3999
$\beta_{3}$ GWR 0.7176 0.2061 0.2091 1.3860
SCC 0.8640 0.5660 0.7871 2.0479
FLAT 0.8718 0.5892 0.8066 2.1975
4 Real data analysis
We apply the FLAT method to study the seawater temperature-salinity (T-S) relationship, which is a critical oceanographic topic with both scientific and practical relevance. In physical ocean dynamics, seawater density is jointly influenced by temperature (which decreases density) and salinity (which increases density). When horizontal gradients of temperature and salinity are spatially opposed, their effects on density become reinforcing (Flament et al., 1985; Rudnick and Luyten, 1996). This phenomenon, known as Temperature-Salinity (T-S) compensation, arises from the counteracting contributions of thermal and haline gradients. Globally, density distributions are largely governed by thermal gradients due to their generally larger magnitude compared to salinity gradients (Reid, 1969; Roden, 1975).
Ocean stratification is fundamentally tied to T-S gradients. Warmer or fresher (less dense) waters tend to overlie colder or saltier (denser) layers, exemplified by Antarctic Intermediate Water. This stratification shapes vertical mixing, nutrient fluxes, and biogeochemical cycles. To quantify spatial gradients, we adopt a numerical method called the spatial difference quotient, which measures the rate of change across discrete spatial datasets-analogous to the derivative in continuous calculus. We introduce these concepts and corresponding formulas in C.
4.1 Dataset and description
We use the public oceanic data from the Copernicus Marine Environment Monitoring Service (CMEMS) [^1] https://marine.copernicus.eu in real data analysis and investigate the T-S relationships via FLAT. CMEMS is a major global ocean observation and forecasting initiative, initiated and coordinated by the European Union, with operational services delivered by Mercator Ocean International (France) to users worldwide. This study utilizes the gridded, daily-averaged physical oceanographic dataset “GLOBAL_MULTIYEAR_001_030” from the GLORYS12V1 product suite under CMEMS. The dataset provides vertical profiles of temperature, salinity, ocean currents, sea level, mixed layer depth, and sea ice parameters, with temporal coverage from 1993 to the present.
This study selected two datasets as the focus of our investigation. The first dataset comprises sea surface T-S data for the North Atlantic Ocean, spanning a meridional range from 0° to 60°W and a zonal range from 0° to 60°N. The second dataset includes T-S observations along a longitudinal transect at 25°W across the Atlantic Basin, extending from 60°N to 60°S latitude and encompassing 43 discrete depth levels between 0 and 3000 meters. This cross-sectional profile, referred to as an oceanographic section, represents a widely adopted approach in marine research for characterizing vertical and horizontal gradients in hydrographic properties.
4.2 Estimation results
We consider the spatial regression modeling between temperature and salinity as
$\mathcal{S}(s_{i})=\beta_{0}(s_{i})+\beta_{1}(s_{i})\mathcal{T}(s_{i})+\epsilon(s_{i}),$ (13)
where $\mathcal{S}$ and $\mathcal{T}$ represent the salinity and temperature of seawater, respectively. Then we optimize the FLAT objective function as below
$L(\bm{\beta})=$ $\frac{1}{n}\sum\limits_{i=1}^{n}\left(\mathcal{S}(s_{i})-\beta_{1}(s_{i})\mathcal{T}(s_{i})-\beta_{0}(s_{i})\right)^{2}$
$+\lambda_{1}\sum\limits_{(s_{i},s_{j})\in\mathbb{E}}\|\beta_{1}(s_{i})-\beta_{1}(s_{j})\|_{1}+\lambda_{2}\sum\limits_{i=1}^{n}\|\beta_{1}(s_{i})\|_{1}$ (14)
and obtain the estimated coefficients for all the spatial locations in the marine data. The tuning and clustering procedures are the same as simulations in Section 3. The evaluation of estimations based on RMSE and the performance of clustering are summarized in Table 3. We find that among the sea level data and ocean section profile data, FLAT exhibits the lowest RMSE, indicating that it provides the best fitting performance. Also, FLAT achieves best clustering results in terms of both SC and CHI in the marine surface data, suggesting an outperforming estimation compared to GWR and SCC. Overall, best results in both datasets demonstrates the advantage of FLAT in detecting the discrete spatial heterogeneity in T-S relationships.
Table 3: Estimation results for T-S relationships in oceanic data
  Method Sea level Sea cross-section
7
13
RMSE SC CHI(×$\times 10^{3}$) RMSE SC CHI(×$\times 10^{3}$)
3
5
7
9
11
13
GWR 0.2879 0.0175 0.4079 0.0827 0.0078 0.1191
SCC 0.1989 0.3293 0.7108 0.0546 0.1093 1.6032
FLAT 0.1941 0.4334 1.9654 0.0500 0.2090 1.8590
4.3 Analysis of the T-S relationship at sea surface layers
In this section, we closely show the estimation results by FLAT on sea-level data in the meridional range from 0° to 60°W and in the zonal range from 0° to 60°N and several valuable findings based on the discrete spatial heterogeneity detected.
The real-world geographical environment of the corresponding area is depicted in Figure 5. Based on the results estimated temperature-salinity relationship coefficients by FLAT in Figure 5, coastal sea areas exhibit significant variations in T-S coefficients due to continental influences. For instance, regions near the eastern Atlantic coast of North America (the top-left region in Figure 5), northeastern South America (the bottom-left region in Figure 5), western Europe (the top-right region ③ in Figure 5), and northwestern Africa (the bottom-left region ④ in Figure 5), particularly estuarine zones, demonstrate pronounced changes in T-S relationships. This phenomenon may arise from localized processes such as evaporation, precipitation, and water mass mixing in coastal waters, including bays and straits, where temperature and salinity variations compensate for each other, thereby affecting seawater density structures.
Additionally, in the open surface waters of the Atlantic Ocean, spanning from 60°N to the equator, T-S coefficients exhibit a pattern of initial decline, followed by a rise and subsequent decrease. This non-monotonic trend may reflect the combined effects of gyral circulation, large-scale current dynamics, and thermohaline processes. In polar-adjacent regions, low temperatures and seasonal freshwater fluxes contribute to distinct salinity patterns, in contrast to equatorial zones, where temperature increases are often accompanied by salinity shifts. These interacting factors jointly shape the spatial variability in T-S relationships. The findings enhance our understanding of thermohaline structures and their variability, which has important implications for the sustainable management of marine systems.
We further calculated the Spatial Difference Quotient (SDQ) of the sea surface T-S relationship based on the estimated coefficients from Figure 5, as shown in Figure 5. The SDQ can be interpreted as a numerical gradient of the T-S relationship at each spatial location; more details on its definition are provided in C. As shown in Figure 5, FLAT effectively captures the spatial variation patterns of T-S relationships. Regions with larger difference quotients are primarily located near landmasses and the equator, particularly near two estuarine zones-the eastern Atlantic coast of North America and northeastern South America. In the equatorial zone, especially at 5°N latitude and between 20°E and 50°E longitude, the T-S coefficient approaches zero, while the SDQ remains relatively high. This pattern likely arises from the distinct marine conditions in equatorial waters: higher temperatures result in reduced density, while intense solar radiation drives strong evaporation and elevated salinity. The counteracting effects of temperature and salinity yield a near-zero T-S coefficient, and the sharp local contrasts contribute to the observed high spatial gradients, reflecting a dynamic T-S compensation regime.
Figure 2: Map of the North Atlantic and its coastal regions
Figure 3: Estimated spatial coefficients for the T-S relationship based on the FLAT model
Figure 4: Estimated local SDQ in T-S coefficients
Figure 5: Identified T-S coefficient contours with clustering
In order to exploring the variation trend of the T-S relationship in the northern Atlantic Ocean more accurately, the spatial clustering of the T-S coefficient is conducted and the detected surface of discrete spatial heterogeneity is shown in Figure 5. The T-S coefficient $\beta_{1}$ is consistently negative ($\beta_{1}<0$) in coastal estuarine regions, indicating a negative correlation between temperature and salinity in these zones. As one moves offshore into the open ocean, $\beta_{1}$ gradually increases and eventually becomes positive ($\beta_{1}>0$), reflecting a positive T-S correlation. Near the equator, contour lines where $\beta_{1}$ approaches zero emerge, forming what is referred to as the inverse temperature-salinity surface. This pattern highlights the spatial transition of T-S relationships across the region and suggests that temperature and salinity changes tend to be balanced in equatorial waters, with the coefficient converging to zero-indicative of a T-S compensation regime. Equatorial surface water, characterized by high temperatures and relatively low salinity, is influenced by intense year-round solar radiation that promotes strong evaporation. However, abundant precipitation delivers substantial freshwater input, offsetting salinity increases and giving rise to a T-S compensation effect, whereby changes in temperature and salinity are counterbalanced and $\beta_{1}$ approaches zero.
4.4 Analysis of the T-S relationship in ocean cross-sections
We also closely analyze the estimated T-S relationship in Atlantic Ocean cross-sections at a longitude of 25°W, spanning a latitude range from 60°N to 60°S, and a depth range from sea level to 3000 meters below sea level.
Figure 8 presents the estimation results of the T-S relationship coefficient $\beta_{1}$ for the ocean section obtained by FLAT. At a depth of approximately 1000 meters, spanning from 60°S latitude to near the equator, there is a notable abrupt change in the T-S relationship, with $\beta_{1}$ transitioning from negative to positive values. This pattern suggests a shift from a negative to a positive correlation between seawater temperature and salinity, which may be attributed to the physical characteristics and circulation patterns of distinct water masses.
In the surface layers of Antarctic waters, temperature is extremely low while salinity is relatively high-a signature consistent with the thermohaline properties of Antarctic Intermediate Water. This is primarily influenced by the cold Antarctic climate, which leads to reduced seawater temperature. Meanwhile, the brine rejection process during sea ice formation increases salinity, thus resulting in a negative T-S correlation.
To better investigate the rate of change in the T-S relationship near the transition zone and more clearly identify thermohaline fronts, we calculated the SDQ of the T-S coefficient, as shown in Figure 8. This metric acts as a discrete analog to the spatial gradient and exhibits a sharp increase near front boundaries, indicating a rapid local change in T-S structure.
To further reveal the spatial patterns of T-S characteristics, we employ the FLAT method to first estimate local coefficients and then conduct clustering to obtain discrete regions with similar thermohaline behavior. Figure 8 displays the contour plot of T-S coefficients across depth and latitude. The results show that the FLAT model successfully delineates T-S regimes in deep ocean layers, capturing the thermohaline features of water masses. Notably, $\beta_{1}$ tends to increase with depth, suggesting that salinity variations contribute more significantly to density changes at greater depths than temperature does.
When $\beta_{1}<0$, temperature and salinity exhibit a negative correlation. Consequently, as seawater temperature decreases, its density increases, simultaneously, a rise in salinity also leads to an increase in density, making the water mass denser and causing the seawater to sink, forming a specific water mass. This process has a significant impact on the global thermohaline circulation. Particularly in deep ocean waters, this phenomenon of density increase facilitates vertical mixing and heat transfer within the ocean. Conversely, when $\beta_{1}>0$, temperature and salinity display a positive correlation. In this scenario, changes in seawater density are influenced by various factors including temperature gradients, salinity gradients, thermal expansion coefficients, and haline contraction coefficients, rendering the situation more complex. Among these factors, if the effects of temperature and salinity on seawater density offset each other, it results in thermohaline compensation.
Each water mass possesses specific temperature and salinity ranges, along with a unique thermohaline relationship that arises from these parameters. For instance, Antarctic Intermediate Water is characterized by low temperature and high salinity, whereas Equatorial Surface Water exhibits high temperature and low salinity. By analyzing temperature and salinity data of water bodies, water mass characteristics can be identified. Furthermore, the thermohaline relationship can also reveal the evolution of water masses. As water masses move and disperse in the ocean, both temperature and salinity undergo changes, reflecting the interaction and evolution process between the water mass and its surrounding environment. Consequently, by analyzing thermohaline data from different locations and times, it is possible to trace the movement trajectory and evolutionary trends of water masses.
Figure 6: Vertical cross-section of the estimated T-S slope coefficients using FLAT
Figure 7: Vertical SDQ of the T-S coefficients along the ocean transect
Figure 8: Identified T-S coefficient contours with clustering in the vertical cross-section
Spatial clustering analysis facilitates a deeper understanding of the variation patterns of the T-S relationship in different sea areas of the Atlantic Ocean and holds significance for marine climate change studies. For instance, water masses such as Antarctic Intermediate Water and equatorial surface water are vital components of the global ocean circulation. Their T-S compensation phenomena influence the operation of the global thermohaline circulation. Warm, low-salinity seawater in equatorial regions is transported to high-latitude areas via ocean currents, where it cools and increases in salinity before sinking to form deep seawater. This process participates in the global oceanic conveyor belt and subsequently exerts significant impacts on global climate.
5 Conclusion and Discussion
We proposed a new approach for modeling spatial heterogeneity, named the Fused Lasso regression model with an Adaptive minimum spanning Tree (FLAT). FLAT is capable of estimating continuous spatial heterogeneity in spatial regression models and simultaneously selecting relevant high-dimensional features, while preserving data-driven detection of discrete spatial heterogeneity through adaptive fused lasso regularization terms. Comprehensive simulation studies demonstrate the excellent finite-sample performance of FLAT. Moreover, the application to oceanographic data on temperature-salinity (T-S) relationships further illustrates the practical usefulness of the proposed method in empirical analyses. While the proposed FLAT framework effectively reconciles continuous and discrete spatial heterogeneity in regression modeling, several potential extensions can be valuable further topics, such as incorporating temporal dynamics into the adaptive fused regularization framework to model spatio-temporal heterogeneity and including uncertainty quantification to improve the interpretability and robustness of the algorithms.
Appendix A The evaluations of clustering
The Rand Index is a crucial metric used to measure the similarity between clustering results and true class labels. If $a$ represents the number of pairs of samples that belong to the same class in both the clustering result and the true class labels, and $b$ represents the number of pairs of samples that belong to different classes in both, the formula for calculating RI is written as
$\text{RI}=\frac{a+b}{\binom{n}{2}}.$ (15)
[…]
$\text{ARI}=\frac{\text{RI}-\text{E}(\text{RI})}{\text{max}(\text{RI})-\text{E}(\text{RI})}.$ (16)
The Silhouette Coefficient (SC) measures clustering effectiveness by evaluating the clustering quality of individual sample points within a cluster. This coefficient comprehensively considers the internal consistency and external separation of sample clusters. The SC ranges from -1 to 1, where a value closer to 1 indicates better clustering performance, a value close to -1 suggests that the sample point is misclustered, and a value near 0 implies uncertainty in determining which cluster a sample belongs to. Assuming there is a sample point $s_{i}$, the average distance between this sample point and samples within the same cluster is denoted as $a(s_{i})$, and the average distance to samples from other clusters is denoted as $b(s_{i})$. Then, the SC for sample point $s_{i}$ is calculated as
$\text{SC}(s_{i})=\frac{b(s_{i})-a(s_{i})}{\max\{a(s_{i}),b(s_{i})\}}.$ (17)
Accordingly, SC for clustering is defined as the average of the SC for all sample points
$\text{SC}=\frac{1}{n}\sum\limits_{i=1}^{n}\text{SC}(s_{i}).$ (18)
Calinski-Harabasz Index (CHI), also called Variance Ratio Criterion (VRC), measures clustering effectiveness using the ratio of the distance between groups and the distance within group. Larger value of CHI indicates better result of clustering. Assuming there are $k$ clusters in the clustering of variable $\mathbf{x}$, with $n_{i}$ samples in each cluster, we note the centroid of cluster $C_{i}$ as $\mathbf{c_{i}}$ and the overall centroid of the data as $\mathbf{c}$. The CHI can be expressed as
$\text{CHI}=\frac{\sum\limits_{i=1}^{k}n_{i}\|\mathbf{c}_{i}-\mathbf{c}\|^{2}/(k-1)}{\sum\limits_{i=1}^{k}\sum\limits_{\mathbf{x}\in C_{i}}\|\mathbf{x}-\mathbf{c}_{i}\|^{2}/(n-k)}.$ (19)
Appendix B Additional numerical results
We show the comparison of different estimations based on all 5 data-generating process in Table 4 below, where the results for DGP1 and DGP5 are also shown also in Table 2 in Section 3.2.
Table 4: Evaluations based on clustering results of estimated parameters $\hat{\bm{\beta}}$ by different methods
Evaluation Dimension Method RI ARI SC CHI(×$\times 10^{4}$)
DGP1 $\beta_{1}$ GWR 0.6411 0.2453 0.4537 0.2041
SCC 0.9117 0.7898 0.8685 2.3816
FLAT 0.9180 0.8051 0.8791 3.3122
$\beta_{2}$ GWR 0.7446 0.2769 0.4163 0.5611
SCC 0.9185 0.7490 0.8378 3.1291
FLAT 0.9289 0.7804 0.8577 4.0711
$\beta_{3}$ GWR 0.8084 0.3576 0.5163 0.9952
SCC 0.9346 0.7709 0.8713 4.3364
FLAT 0.9370 0.7818 0.8835 4.7318
DGP2 $\beta_{1}$ GWR 0.6220 0.1843 0.3160 0.2232
SCC 0.8816 0.7280 0.8373 1.5141
FLAT 0.8932 0.7519 0.8492 2.0688
$\beta_{2}$ GWR 0.7210 0.2173 0.3331 0.5846
SCC 0.8942 0.6826 0.8137 2.4471
FLAT 0.9074 0.7176 0.8307 3.2222
$\beta_{3}$ GWR 0.7800 0.2933 0.4142 1.0047
SCC 0.9002 0.6753 0.8432 2.9139
FLAT 0.9107 0.7056 0.8441 3.6620
DGP3 $\beta_{1}$ GWR 0.6298 0.1702 0.2606 0.2975
SCC 0.8525 0.6607 0.7910 1.2801
FLAT 0.8697 0.6946 0.8236 1.8807
$\beta_{2}$ GWR 0.7027 0.1853 0.2720 0.6311
SCC 0.8732 0.6185 0.7928 2.0624
FLAT 0.8904 0.6672 0.8204 2.6647
$\beta_{3}$ GWR 0.7619 0.2607 0.3454 1.0801
SCC 0.8832 0.6296 0.8217 2.4289
FLAT 0.8959 0.6589 0.8281 2.8204
DGP4 $\beta_{1}$ GWR 0.6206 0.1481 0.2246 0.3315
SCC 0.8286 0.6071 0.7702 1.0753
FLAT 0.8516 0.6545 0.7990 1.5322
$\beta_{2}$ GWR 0.6968 0.1681 0.2144 0.7392
SCC 0.8559 0.5701 0.7715 1.8524
FLAT 0.8674 0.6030 0.7841 2.5714
$\beta_{3}$ GWR 0.7467 0.2370 0.2862 1.1285
SCC 0.8661 0.5871 0.7931 2.0760
FLAT 0.8799 0.6150 0.8172 2.4064
DGP5 $\beta_{1}$ GWR 0.6101 0.1260 0.1879 0.3858
SCC 0.8056 0.5650 0.7494 0.7449
FLAT 0.8227 0.5951 0.7680 1.0421
$\beta_{2}$ GWR 0.6760 0.1487 0.1363 0.7457
SCC 0.8307 0.5023 0.7572 1.6862
FLAT 0.8471 0.5386 0.7754 2.3999
$\beta_{3}$ GWR 0.7176 0.2061 0.2091 1.3860
SCC 0.8640 0.5660 0.7871 2.0479
FLAT 0.8718 0.5892 0.8066 2.1975
Appendix C Details of the real-data analysis
Taking two-dimensional space as an example, suppose an observation point $s_{i}$ , whose gradient points projected onto the horizontal and vertical axes are denoted as $s_{i_{1}}$ and $s_{i_{2}}$, respectively. If the observed value at each point $s$ in space is represented by $\beta(s)$, then the Spatial Difference Quotient (SDQ) at point $s_{i}$ can be expressed as:
$D(s_{i})=\sqrt{\frac{\left(\beta(s_{i})-\beta(s_{i_{1}})\right)^{2}}{d_{1}^{2}}+\frac{\left(\beta(s_{i})-\beta(s_{i_{2}})\right)^{2}}{d_{2}^{2}}},$ (20)
Here, $d_{1}$ and $d_{2}$ denote the Euclidean distances from point $s_{i}$ to $s_{i_{1}}$ and $s_{i_{2}}$, respectively. Since the observed values at the two gradient points in the aforementioned formula may not exist, we instead consider the two nearest observation points to $s_{i}$, denoted as $s_{j}$ and $s_{k}$. Consequently, $D(s_{i})$ is defined as:
$D(s_{i})=\sqrt{\frac{\left(\beta(s_{i})-\beta(s_{j})\right)^{2}}{d_{ij}^{2}\sin^{2}\gamma}+\frac{\left(\beta(s_{i})-\beta(s_{k})\right)^{2}}{d_{ik}^{2}\sin^{2}\gamma}-2\frac{\left(\beta(s_{i})-\beta(s_{j})\right)\left(\beta(s_{i})-\beta(s_{k})\right)\cos\gamma}{d_{ij}d_{ik}\sin^{2}\gamma}},$ (21)
Here, $\gamma$ denotes the included angle between vector $\overrightarrow{s_{i}s_{j}}$ and vector $\overrightarrow{s_{i}s_{k}}$.